OpenStudy (jhulian):

Find the limit of s(n) as n → ∞. s(n) = (1/n^2)*[n(n + 1)/8]

OpenStudy (irishboy123):

is it \(s(n) = (\dfrac{1}{n^2})\dfrac{n(n + 1)}{8}\) ?? if not, maybe draw it??

OpenStudy (jhulian):

sorry, that is it

OpenStudy (jhulian):

\[s(n) = \frac{ 1 }{ n^2 }\left[ \frac{ n(n+1) }{ 8} \right]\]

OpenStudy (irishboy123):

divide it out! \[s(n) = \frac{ 1 }{ n^2 }\left[ \frac{ n^2+n }{ 8} \right]\] \[s(n) =\left[ \frac{ 1+ \frac{1}{n} }{ 8} \right]\]

OpenStudy (irishboy123):

see it?

OpenStudy (jhulian):

oh yeah...

OpenStudy (irishboy123):

so if \(n \to \infty\) , then get \(n\) underneath something so **the something** goes to zero......

OpenStudy (jhulian):

um... so what is the final answer?

OpenStudy (daniel.ohearn1):

You may notice that IrishBoy divided each term by the largest value of n, n^2, getting another form of the same function. Now what is the limit as n approaches infinity of 1/n ?

OpenStudy (jhulian):

zero?

OpenStudy (daniel.ohearn1):

Yes, so what must the limit be?

OpenStudy (jhulian):

1/8

OpenStudy (daniel.ohearn1):

Right.

OpenStudy (jhulian):

Thank you Daniel

OpenStudy (daniel.ohearn1):

np